OpenStudy (anonymous):

The path of a soccer ball is modeled by the function h(x) = -0.005x2 + 0.25x, where h is the height in meters and x is the horizontal distance that the ball travels in meters. What is the maximum height that the ball reaches?

OpenStudy (y2o2):

get the vertex Vx - -b/2a = -0.25/(2*-0.005) = 25 Vy = -0.005(25)² + 0.25(25)= 3.125 meters